# GRAVITY: A Framework for Personalized Text Generation via Profile-Grounded Synthetic Preferences

### v1

@@ 1 Introduction
Personalization has become a critical frontier for LLMs Jang et al. (2023); Chen et al. (2024); Zhang et al. (2024b). While recent advances enable fluent and contextually relevant text generation, outputs often remain generic, overlooking individual differences in taste, style, and preference Zhang et al. (2025); Moorjani et al. (2022). In domains such as book recommendations, this gap is particularly visible—two users may value entirely different aspects of the same content, with one drawn to narrative structure and cultural context and the other to character development and personal resonance. Generic descriptions risk alienating users by failing to capture what truly matters Yunusov et al. (2024); Cai et al. (2023).
Prior work in personalization often relies on human-annotated preference data, whether through reinforcement learning from human feedback (RLHF) Kirk et al. (2024); Poddar et al. (2024), preference modeling Lee et al. (2024a); Zhong et al. (2024), or profile-conditioned prompting Zhang et al. (2024a); Lyu et al. (2023). While effective, large-scale annotation is costly and difficult to scale, limiting the breadth of user attributes that can be incorporated. Moreover, existing approaches frequently reduce personalization to narrow signals such as demographics or explicit traits, overlooking deeper dimensions of user engagement.
To combat this challenge, we introduce GRAVITY (Generative Response with Aligned Values, Interests, and Traits of You), a framework for creating synthetic, profile-grounded preference data that can be used to fine-tune LLMs for more effective personalization. As a case study, we focus on personalizing book descriptions using users from the Amazon Book Reviews dataset Hou et al. (2024). Instead of collecting explicit preference annotations, we construct synthetic preference pairs grounded in well-established psychological and cultural frameworks—including Hofstede’s cultural dimensions Hofstede (1983), Schwartz’s theory of basic values Schwartz (2012), the World Values Survey Haerpfer et al. (2024), and the Big Five OCEAN traits Goldberg (2013). This allows us to capture variation in user interests, values, beliefs, and personality traits, providing a richer basis for personalization than demographics alone. We then fine-tune Llama-3.1-8B-Instruct with Direct Preference Optimization (DPO) Rafailov et al. (2023), aligning generations with these profile-derived preferences.
@@ 2 Background
Personalization is central to making generated content engaging and relevant: attributes such as user interests, style preferences (especially through personality) Nguyen et al. (2018), or cultural background Matz et al. (2024) can strongly shape whether stories and characters resonate with readers Woźniak et al. (2024); Yang et al. (2023). Recent work leverages LLMs for personalization across both direct text generation and downstream applications like recommendation Jiang et al. (2023); Jang et al. (2023). Broadly, methods fall into two streams: (a) personalized generation, where models adapt outputs via prompting, user profiles, or fine-tuning Peng et al. (2024); Li et al. (2024b); and (b) LLMs for personalized tasks, such as recommendation, reasoning over user histories, or tailoring explanations Bismay et al. (2024); Shao et al. (2024). Hybrid approaches increasingly combine these, generating user-specific content that simultaneously improves alignment and task performance Shenfeld et al. (2025); Zhong et al. (2024).
Within generation, personalization extends beyond static profile conditioning to new task formulations. Work on summarization, for example, uses multi-agent pipelines to refine drafts according to user preferences Xiao et al. (2023), while headline generation has incorporated user-history modeling Song et al. (2023). Similarly, recommender systems use LLMs to expand sparse item text or construct structured user/item profiles from histories, enabling personalized reviews and explanations Lyu et al. (2023); Acharya et al. (2023).
Finally, advances in alignment and evaluation point toward more fine-grained personalization. Extensions of RLHF introduce per-user reward components or factorized preference functions Li et al. (2024b); Shenfeld et al. (2025); Poddar et al. (2024), while controllable preference vectors allow multi-objective adaptation Zhong et al. (2024). Benchmarks like LaMP Salemi et al. (2023) and newer evaluations on role-playing, user modeling, and cultural adaptation Tseng et al. (2024); Zhang et al. (2024b) reflect this growing focus on personalization as a key axis of LLM evaluation.
@@ 3.2 Creating User Profiles
Each user profile in our framework consists of four key components: (1) Demographics (gender, age, location), (2) Interests, (3) Values and Beliefs, (4) Personality Traits. Demographic information is obtained directly from user-provided data when available. When gender or age is missing, we follow prior work that has shown strong performance in demographic estimation and apply trained DeBERTa models Tureyyen (2023a, b). To reduce noise and sparsity, we bin ages into three groups: young (<30), middle-aged (31–60), and senior (>60).
@@
Personality Traitscapture a user’s underlying preferences and tendencies, shaping both their writing style and the kinds of story elements they are likely to connect most with. To infer these traits, we use PersonalityLM Wang and Sun (2024), a fine-tuned RoBERTa classifier to predict high/low levels for each of the Big Five (OCEAN) traits. In our study, we estimate user personalities using each user’s historical book reviews.
@@ 3.3 Generating Custom Preferences
Based on users’ generated profiles, the next step in GRAVITY is to construct custom preference data that guides model adaptation toward personalized product descriptions. Rather than relying on generic preference statements, we generate scenario-based data that captures how a user’s interests, values and beliefs, and personality traits shape their judgments in concrete contexts. While the underlying pool of scenarios remains consistent across users, each profile produces a unique mapping of chosen and rejected labels, ensuring that supervision reflects their individual preferences. This scenario-driven approach provides richer, more discriminative signals than abstract descriptions, enabling the model to learn finer-grained distinctions in how different users engage with content Singh et al. (2025); Huang et al. (2023).
@@
Personality TraitsWe leverage two situation-based personality questionnaires, TRAIT (scenarios) Lee et al. (2024b) and Big5Chat (dialogues) Li et al. (2024a), each containing questions with binary-level answers for the OCEAN traits. For each user, we randomly select 150 question–answer pairs from each dataset. For each pair, the answer consistent with the user’s OCEAN trait profile is labeled chosen, and the alternative labeled rejected.
@@ 3.4 User-Centric LLM Training
For our task of personalized book descriptions, we generate approximately 1,000 customized preference pairs across all three dimensions for each user (400K for user base). Using this data, we utilize Direct Preference Optimization (DPO) Rafailov et al. (2023) to generate a single personalized LLM which learns the behaviors and styles of the users in our dataset. Each training instance conditions the model on a user’s demographic, personality, values, and interests, enabling generalization across diverse profiles. At inference, we provide the target user’s profile to generate personalized book descriptions. For evaluation, we sample a highly rated title from the user’s preferred genres, excluding books they have already reviewed. Additional training details are in §4.1.
@@ 3.5 Product Description Personalization
To generate personalized descriptions, we follow prior work Santurkar et al. (2023); Hwang et al. (2023) by conditioning model responses on the user’s profile. Each profile, as described in §3.2, includes demographics, genre interests, personality traits, and personal values. Full prompt templates for description generation are provided in §C. For our task, we select a target book description to personalize from a randomly chosen highly rated title within the user’s top genres, ensuring the user has not previously purchased or reviewed the book.
@@ 4.1 Experiment Setup
Finetuning user-centric LLMs We adapt Llama-3.1-8B-Instruct Dubey et al. (2024) to user preferences using DPO with LoRA adapters Hu et al. (2022). We preference tune the model using the TRL library CarperAI (2023) with the AdamW optimizer Loshchilov and Hutter (2017), a $\beta$ value of 0.3, a learning rate of $2\times 10^{-5}$, weight decay of 0.01, a cosine learning rate schedule, and a warmup ratio of 0.05. Additional training details and prompt formats are detailed in §D.1.
Personalization strategies We benchmark against seven baselines [^4] All methods use Llama-3.1-8B-Instruct unless noted., differing in how user information is integrated: (1) Prompting: (a) BaseRewrite — rewrites the description to be more engaging (no personalization), (b) DemoBased — appends user demographics (age, gender, location), (c) UserSummary — prepends a GPT-4o–generated user profile summary, (d) LaMP Salemi et al. (2023) — retrieves and prepends the user’s most relevant past reviews, (e) TriAgent Xiao et al. (2023) — a three-stage pipeline combining BaseRewrite, GPT-4o summaries, and customized edit instructions; (2) SFT: (a) UserSFT — fine-tunes on (book description → user review) pairs with persona prompts encoding user demographics; (3) Preference data: (a) PrefAlign — uses DPO using GPT-4o-generated aligned and misaligned book descriptions, produced solely from user demographics. Details on baseline implementations are in §D.2.
Auto-evaluation metrics To assess the quality of personalized descriptions, we utilize three personalization-measuring metrics: (1) Top-1 WinRate: how often is a method’s output preferred over all other methods, (2) Preference Gain: how often is a method’s output preferred over the original description, and (3) Interestingness Score: how engaging/captivating does the user find the method’s output (based on 5-point Likert score). We evaluate personalizations using LLM-as-a-judge Zheng et al. (2023). We provide GPT-4o with a detailed user persona, generated from the user’s complete set of written reviews and and available demographics. To ensure personalized content remains faithful to the original description, we also compute cosine similarity between SentenceBERT Reimers and Gurevych (2019) embeddings of the original and generated texts. Prompt templates used for evaluation are included in §D.3.
@@ 4.2 Personalization with GRAVITY
Personalization Across Genres We further analyze how personalization effectiveness varies across book genres. While our method consistently improves over baselines across both fiction and nonfiction, we observe substantially higher gains for fiction titles (see Figure 3555Although the Amazon Reviews dataset has several fine-grained labels for book genres, in this analysis, we focus on a set of nine hierarchical book categories/themes.); we see preference gainst above 75% for fiction books but gains closer to 65% for non-fiction books. Our findings align with prior research which show that fiction readers tend to be more engaged with narrative elements—such as character development, emotional arcs, and imaginative settings—which naturally connect with readers’ personal values, beliefs, and personality traits Bal and Veltkamp (2013); Goyal and Mahmoud (2024). Thus, incorporating these data as training signals into user-centric models allows for better personalization. In contrast, nonfiction reading is primarily fact-oriented, emphasizing clarity, accuracy, and logical structure, which limits the potential for personalization, resulting in smaller improvements through GRAVITY.
@@ Appendix B Custom Preference Generation
Personality TraitsTo generate personality preference pairs, we utilize two existing datasets: TRAIT Lee et al. (2024b) and Big5Chat Li et al. (2024a). In Table 10, we summarize main statistics from these two datasets. We randomly sample 300 total questions (60 questions for each OCEAN trait, 30 from each dataset) [^6] As TRAIT contains 2 answers for each binary label answer, we randomly select one of the answers from each level to generate a preference pair Table 11 contains example questions from each dataset.
@@
TRAIT Lee et al. (2024b) 5,000 (400 per trait) Situation-based Binary (2 answers per low/high)
Big5Chat Li et al. (2024a) 100,000 (20K per trait) Dialogue-based Binary
@@ D.1 Finetuning User-Centric LLMs
To finetune Llama-3.2-3B-Instruct, we use DPO and 4-bit quantized LoRA. We preference-tune a single model for all users using the TRL library. We use a final learning rate of: $2\times 10^{-5}$ and $\beta$ value of: 0.3. We use a batch size of 64 with a gradient accumulation step size of 2 and 3 epochs with early stopping criterion. We train our model on 2 A6000s with DeepSpeed Rajbhandari et al. (2020) which takes approximately 11 hours. Table D.2 contains our training prompts for preference aligning the model.
@@ D.2 Personalized Prompting Strategies
For the LaMP-based approach, we first identify the most relevant book descriptions to the target book recommendation. To do this, we first encode all the descriptions of the books a user has reviewed into dense vector representations using a SentenceBERT (all-MiniLM-L6-v2) model. Each description is thus mapped to a high-dimensional embedding capturing its semantic content. Given a target book, we similarly encode its description into the same embedding space and retrieve the top-5 most similar books based on cosine similarity. For efficient similarity search, we leverage FAISS Douze et al. (2024) to index and query the embeddings. Based on the extracted similar books, we prepend the user’s reviews for these books as context in the generation prompt.
@@ D.4 User Study Design
We further measure users’ personality traits using the Mini-IPIP questionnaire Donnellan et al. (2006), a validated 20-item Likert-scale instrument for assessing the Big Five (OCEAN) traits. To estimate users’ values and beliefs, we adapt our 150-item seed statement bank, selecting a concise subset of 10 statements, across various categories: culture, ethics, society, politics, morals, and religion, to minimize participant fatigue. We then prompt GPT-4o to infer users’ likely responses to the remaining statements based on their annotated subset:

### v2

@@ 1 Introduction
Personalization has become a critical frontier for LLMs jang2023personalized; chen2024persona; zhang2024personalization. While recent advances enable fluent and contextually relevant text generation, outputs often remain generic, overlooking individual differences in taste, style, and preference zhang-etal-2025-personalized; moorjani-etal-2022-audience. In domains such as book recommendations, this gap is particularly visible—two users may value entirely different aspects of the same content, with one drawn to narrative structure and cultural context and the other to character development and personal resonance. Generic descriptions risk alienating users by failing to capture what truly matters yunusov2024mirrorstoriesreflectingdiversitypersonalized; cai-etal-2023-generating.
Prior work in personalization often relies on human-annotated preference data, whether through reinforcement learning from human feedback (RLHF) kirk2024prism; poddar2024personalizing, preference modeling lee2024aligning; zhong2024panacea; zheng2025tuningfreellmbuildstrong, or profile-conditioned prompting zhang2024p4; lyu2023llm. While effective, large-scale annotation is costly and difficult to scale, limiting the breadth of user attributes that can be incorporated. Moreover, existing approaches frequently reduce personalization to narrow signals such as demographics or explicit traits, overlooking deeper dimensions of user engagement.
To combat this challenge, we introduce GRAVITY (Generative Response with Aligned Values, Interests, and Traits of You), a framework for creating synthetic, profile-grounded preference data that can be used to fine-tune LLMs for more effective personalization. As a case study, we focus on personalizing book descriptions using users from the Amazon Book Reviews dataset hou2024bridging. Instead of collecting explicit preference annotations, we construct synthetic preference pairs grounded in well-established psychological and cultural frameworks—including Hofstede’s cultural dimensions hofstede1983national, Schwartz’s theory of basic values schwartz2012overview, the World Values Survey Haerpfer2024-jz, and the Big Five OCEAN traits goldberg2013alternative. This allows us to capture variation in user interests, values, beliefs, and personality traits, providing a richer basis for personalization than demographics alone. We then fine-tune Llama-3.1-8B-Instruct with Direct Preference Optimization (DPO) rafailov2023direct, aligning generations with these profile-derived preferences.
@@ 2 Background
Personalization is central to making generated content engaging and relevant: attributes such as user interests, style preferences (especially through personality) nguyen2018user; dey2025can, or cultural background matz2024potential; joshi2025improving; liu2025can can strongly shape whether stories and characters resonate with readers wozniak2024personalized; yang2023palr. Recent work leverages LLMs for personalization across both direct text generation and downstream applications like recommendation jiang2023personallm; jang2023personalized. Broadly, methods fall into two streams: (a) personalized generation, where models adapt outputs via prompting, user profiles, or fine-tuning peng2024review; li2024personalized; and (b) LLMs for personalized tasks, such as recommendation, reasoning over user histories, or tailoring explanations bismay2024reasoningrec; shao2024ulmrec. Hybrid approaches increasingly combine these, generating user-specific content that simultaneously improves alignment and task performance shenfeld2025language; zhong2024panacea.
Within generation, personalization extends beyond static profile conditioning to new task formulations. Work on summarization, for example, uses multi-agent pipelines to refine drafts according to user preferences xiao2023personalized, while headline generation has incorporated user-history modeling song2023general. Similarly, recommender systems use LLMs to expand sparse item text or construct structured user/item profiles from histories, enabling personalized reviews and explanations lyu2023llm; acharya2023llm.
Finally, advances in alignment and evaluation point toward more fine-grained personalization. Extensions of RLHF introduce per-user reward components or factorized preference functions li2024personalized; shenfeld2025language; poddar2024personalizing, while controllable preference vectors allow multi-objective adaptation zhong2024panacea. Benchmarks like LaMP salemi2023lamp and newer evaluations on role-playing, user modeling, and cultural adaptation tseng2024two; zhang2024personalization reflect this growing focus on personalization as a key axis of LLM evaluation.
@@ 3.2 Creating User Profiles
Each user profile in our framework consists of four key components: (1) Demographics (gender, age, location), (2) Interests, (3) Values and Beliefs, (4) Personality Traits. Demographic information is obtained directly from user-provided data when available. When gender or age is missing, we follow prior work that has shown strong performance in demographic estimation and apply trained DeBERTa models kaantureyyen2023deberta_age; kaantureyyen2023deberta_gender. To reduce noise and sparsity, we bin ages into three groups: young (<30), middle-aged (31–60), and senior (>60).
@@
Personality Traitscapture a user’s underlying preferences and tendencies, shaping both their writing style and the kinds of story elements they are likely to connect most with. To infer these traits, we use PersonalityLM wang2024continuous, a fine-tuned RoBERTa classifier to predict high/low levels for each of the Big Five (OCEAN) traits. In our study, we estimate user personalities using each user’s historical book reviews.
@@ 3.3 Generating Custom Preferences
Based on users’ generated profiles, the next step in GRAVITY is to construct custom preference data that guides model adaptation toward personalized product descriptions. Rather than relying on generic preference statements, we generate scenario-based data that captures how a user’s interests, values and beliefs, and personality traits shape their judgments in concrete contexts. While the underlying pool of scenarios remains consistent across users, each profile produces a unique mapping of chosen and rejected labels, ensuring that supervision reflects their individual preferences. This scenario-driven approach provides richer, more discriminative signals than abstract descriptions, enabling the model to learn finer-grained distinctions in how different users engage with content singh2025fspofewshotpreferenceoptimization; huang2023learning.
@@
Personality TraitsWe leverage two situation-based personality questionnaires, TRAIT (scenarios) Lee2024TRait and Big5Chat (dialogues) li2024big5chat, each containing questions with binary-level answers for the OCEAN traits. For each user, we randomly select 150 question–answer pairs from each dataset. For each pair, the answer consistent with the user’s OCEAN trait profile is labeled chosen, and the alternative labeled rejected.
@@ 3.4 User-Centric LLM Training
For our task of personalized book descriptions, we generate approximately 1,000 customized preference pairs across all three dimensions for each user (400K for user base). Using this data, we utilize Direct Preference Optimization (DPO) rafailov2023direct to generate a single personalized LLM which learns the behaviors and styles of the users in our dataset. Each training instance conditions the model on a user’s demographic, personality, values, and interests, enabling generalization across diverse profiles. At inference, we provide the target user’s profile to generate personalized book descriptions. For evaluation, we sample a highly rated title from the user’s preferred genres, excluding books they have already reviewed. Additional training details are in §4.1.
@@ 3.5 Product Description Personalization
To generate personalized descriptions, we follow prior work santurkar2023whose; hwang2023aligning by conditioning model responses on the user’s profile. Each profile, as described in §3.2, includes demographics, genre interests, personality traits, and personal values. Full prompt templates for description generation are provided in §C. For our task, we select a target book description to personalize from a randomly chosen highly rated title within the user’s top genres, ensuring the user has not previously purchased or reviewed the book.
@@ 4.1 Experiment Setup
Finetuning user-centric LLMs We adapt Llama-3.1-8B-Instruct dubey2024llama to user preferences using DPO with LoRA adapters hu2022lora. We preference tune the model using the TRL library trl2023 with the AdamW optimizer loshchilov2017decoupled, a $\beta$ value of 0.3, a learning rate of $2\times 10^{-5}$, weight decay of 0.01, a cosine learning rate schedule, and a warmup ratio of 0.05. Additional training details and prompt formats are detailed in §D.1.
Personalization strategies We benchmark against seven baselines [^4] All methods use Llama-3.1-8B-Instruct unless noted., differing in how user information is integrated: (1) Prompting: (a) BaseRewrite — rewrites the description to be more engaging (no personalization), (b) DemoBased — appends user demographics (age, gender, location), (c) UserSummary — prepends a GPT-4o–generated user profile summary, (d) LaMP salemi2023lamp — retrieves and prepends the user’s most relevant past reviews, (e) TriAgent xiao2023personalized — a three-stage pipeline combining BaseRewrite, GPT-4o summaries, and customized edit instructions; (2) SFT: (a) UserSFT — fine-tunes on (book description → user review) pairs with persona prompts encoding user demographics; (3) Preference data: (a) PrefAlign — uses DPO using GPT-4o-generated aligned and misaligned book descriptions, produced solely from user demographics. Details on baseline implementations are in §D.2.
Auto-evaluation metrics To assess the quality of personalized descriptions, we utilize three personalization-measuring metrics: (1) Top-1 WinRate: how often is a method’s output preferred over all other methods, (2) Preference Gain: how often is a method’s output preferred over the original description, and (3) Interestingness Score: how engaging/captivating does the user find the method’s output (based on 5-point Likert score). We evaluate personalizations using LLM-as-a-judge zheng2023judgingllmasajudgemtbenchchatbot. We provide GPT-4o with a detailed user persona, generated from the user’s complete set of written reviews and and available demographics. To ensure personalized content remains faithful to the original description, we also compute cosine similarity between SentenceBERT reimers2019sentence embeddings of the original and generated texts. Prompt templates used for evaluation are included in §D.3.
@@ 4.2 Personalization with GRAVITY
Personalization Across Genres We further analyze how personalization effectiveness varies across book genres. While our method consistently improves over baselines across both fiction and nonfiction, we observe substantially higher gains for fiction titles (see Figure 3555Although the Amazon Reviews dataset has several fine-grained labels for book genres, in this analysis, we focus on a set of nine hierarchical book categories/themes.); we see preference gainst above 75% for fiction books but gains closer to 65% for non-fiction books. Our findings align with prior research which show that fiction readers tend to be more engaged with narrative elements—such as character development, emotional arcs, and imaginative settings—which naturally connect with readers’ personal values, beliefs, and personality traits bal2013does; goyal2024systematic. Thus, incorporating these data as training signals into user-centric models allows for better personalization. In contrast, nonfiction reading is primarily fact-oriented, emphasizing clarity, accuracy, and logical structure, which limits the potential for personalization, resulting in smaller improvements through GRAVITY.
@@ Appendix B Custom Preference Generation
Personality TraitsTo generate personality preference pairs, we utilize two existing datasets: TRAIT Lee2024TRait and Big5Chat li2024big5chat. In Table 10, we summarize main statistics from these two datasets. We randomly sample 300 total questions (60 questions for each OCEAN trait, 30 from each dataset) [^6] As TRAIT contains 2 answers for each binary label answer, we randomly select one of the answers from each level to generate a preference pair Table 11 contains example questions from each dataset.
@@
TRAIT Lee2024TRait 5,000 (400 per trait) Situation-based Binary (2 answers per low/high)
Big5Chat li2024big5chat 100,000 (20K per trait) Dialogue-based Binary
@@ D.1 Finetuning User-Centric LLMs
To finetune Llama-3.2-3B-Instruct, we use DPO and 4-bit quantized LoRA. We preference-tune a single model for all users using the TRL library. We use a final learning rate of: $2\times 10^{-5}$ and $\beta$ value of: 0.3. We use a batch size of 64 with a gradient accumulation step size of 2 and 3 epochs with early stopping criterion. We train our model on 2 A6000s with DeepSpeed rajbhandari2020zeromemoryoptimizationstraining which takes approximately 11 hours. Table D.2 contains our training prompts for preference aligning the model.
@@ D.2 Personalized Prompting Strategies
For the LaMP-based approach, we first identify the most relevant book descriptions to the target book recommendation. To do this, we first encode all the descriptions of the books a user has reviewed into dense vector representations using a SentenceBERT (all-MiniLM-L6-v2) model. Each description is thus mapped to a high-dimensional embedding capturing its semantic content. Given a target book, we similarly encode its description into the same embedding space and retrieve the top-5 most similar books based on cosine similarity. For efficient similarity search, we leverage FAISS Douze2024TheFL to index and query the embeddings. Based on the extracted similar books, we prepend the user’s reviews for these books as context in the generation prompt.
@@ D.4 User Study Design
We further measure users’ personality traits using the Mini-IPIP questionnaire donnellan2006mini, a validated 20-item Likert-scale instrument for assessing the Big Five (OCEAN) traits. To estimate users’ values and beliefs, we adapt our 150-item seed statement bank, selecting a concise subset of 10 statements, across various categories: culture, ethics, society, politics, morals, and religion, to minimize participant fatigue. We then prompt GPT-4o to infer users’ likely responses to the remaining statements based on their annotated subset: